OpenStudy (anonymous):

what does log mean? And how would you solve this? y = log x If y = 10, then what is x?

OpenStudy (amistre64):

y=log(x) we assume base 10 if there is no "base" included 10 = log(x) 10^x = 10 x = 1

OpenStudy (amistre64):

or do I have that backwards....

OpenStudy (amistre64):

x = 10,000,000,000 maybe.. 10 = log(x) 10^10 = 10^log(x) 10^10 = x yep... had it backwards

OpenStudy (anonymous):

Thanks Actually i think you were right the first time because i though it out y = log x 10 = log x 10 = 10^x x = 1

OpenStudy (amistre64):

no, I had it wrong first time, so I worked it out with the rules... your third step there, you need to base10 both sides... 10^10 = 10^log(x) the right hand side become x now 10^10 = x x = 10,000,000,000

OpenStudy (amistre64):

another way to look at is is logs ARE exponents, 10 is the exponent of your base 10. 10^10 = x....

OpenStudy (anonymous):

http://library.thinkquest.org/J002235/hard.html and could you give me any webites on the work involved in log?

OpenStudy (amistre64):

log2(x) = 3 2^3 = x x = 8 or.... log2(x) = 3 2^log2(x) = 2^3 x = 2^3 x = 8

OpenStudy (amistre64):

websites? nah.... I just work them out of the library textbooks if anything

OpenStudy (anonymous):

thanks a lot your help is gadly appreciated